HUBBLE SPACE TELESCOPE DAILY REPORT #1210 PERIOD COVERED: 0000Z Daily Status Report as of 237/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5342 (UV Imaging of Intermediate- Age Magellanic Cloud Clusters) The WF/PC-2 was used to make ultraviolet observations of the star cluster NGC-1866 which is located in the Large Magellanic Cloud to determine the nature of its hot stellar components. The observations were completed as planned, and no problems were reported. 1.2 Completed WF/PC-2 5655 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats Part 2) The WF/PC-2 was used to make visual flat calibration images. The observations were completed as planned, and no problems were reported. 1.3 Completed GHRS/2 5598 (GHRS Side 2 Aperture Offsets: Cycle 4) The High Resolution Spectrograph (Side Two) was used to make wave-length calibrations and observations of standard star HD141556 to determine the wave-length offsets between the Large Science Aperture (LSA) and the Small Science Aperture (SSA) for Side Two gratings. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 5476 (Continuum Snapshots of 3CR Radio Galaxies) The WF/PC-2 was used to make snapshot images of radio galaxy 3C454.2 as part of a continuum survey and to study the galaxy's morphology. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed FOS/RD 5638 (High Resolution Observations of Nova CYG 92: Cycle 4) The Faint Object Spectrograph (Red side) will be used to make low resolution spectral observations of Nova Cygni 1992 to determine the metal abundance and physical conditions in the nova ejecta. The observations were completed as planned, and no problems were reported. 1.6 Completed WF/PC-2 5371 (HST Medium_ Deep Survey: Cycle 4 Medium Low Galactic Latitude) The WF/PC-2 was used in parallel with the prime science observations to make low galactic latitude deep multicolor images of objects such as galaxies, and stars in our galaxy. The observations were completed as planned, and no other problems were reported. 1.7 Completed Two Sets of WF/PC-2 5571 (WF/PC-2 Cycle 4 Calibration: Earthflats, Smaller Set) The WF/PC-2 was used to make several ultraviolet and narrow-band filter Earth flat calibration observations to correct for the telescope's illumination pattern. The observations were completed as planned, and no problems were reported. 1.8 Completed Two Sets of WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain several 30-minute long, dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 5642 (Observations of the Impact of Comet Shoemaker-Levy (1993E) and Jupiter) The WF/PC-2 was used to make observations to support the Jupiter/Shoemaker-Levy event. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 12 Successful : 12 2.2 FHST Updates: Scheduled :23 Successful :23 2.3 Significant HSTARs: None 2.4 Operations Notes: No anomaly reports have been filed by any HST subsystem during the past operational day. The Battery #1 Capacity Test was completed on schedule with the battery now fully charged. It is not yet, however, being used in the online string. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-